Shift report 20011027 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Oct 27 2001 - 16:06:36 EDT

  • Next message: Apache: "Shift report 20011028 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Oct 27 16:06:36 2001, Shift: 08:00-16:00
    Supervisor: J.I.Jordre, with: 
    Log book#10 pages 62 - 63
    
    Shift Summary:
    ------------------------------------------------------------
    1. Cryo problems from this morning continued until 12:50. However, some
       Booster main magnet power supply was giving problems and the first 
       attempt to fill didn't occur until 15:10. Then some attemted fills in 
       both rings, with correponding dumps. Finally attempted ramp around
       15:25, but blue beam was lost.
    2. TPC pedestals, pedestals and DVM runs 5535-5537.
    3. Yeah, right!
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Oct 27 2001 - 16:07:14 EDT